### 14 Dec 2023 v3

@@ 5.3 Employee productivity
Data-driven decision-making, or the forecasting of required outputs to estimate the required workforce, is an excellent way to minimize such costs of hiring and laying-off by adequately scheduling the workforce [6]. Through proper scheduling, the workforce from idle time can be shifted for workdays requiring extra hours and thereby balanced. [100] showed that work pressure could be balanced with reduced slack time and workforce through different heuristic algorithms, with each algorithm performing well in different areas of efficiency.
@@ 6.1 Supply chain performance
Immediately after the tasks are completed, the performance data must be recorded. For data collection, the performance metrics are first to be identified, just as those found for business evaluations [15]. [51] mentioned plan success, source optimization, production efficiency, delivery performance, and customer support-relation and satisfaction, each having multiple performance metrics under them. The performance level found afterward can be of three types below average, average, and above-average [5]. The actions followed after such a finding is different in each case. When a below-average performance is observed, managers can either look for anomalies in the system or review whether the goals set were too high to achieve.
@@ 6.2.5 Mean Absolute Scaled Error
With MAE, outliers are protected; with RMSE, we are assured of an impartial prediction. SC Analysts need to analyze MAE and see whether it results in a significant bias, and therefore they should utilize RMSE. In situations when there are many outliers in the dataset, MAE may help correct the skewed prediction.
@@ 6.2.6 Tracking Signal
The tracking signal is the way to verify if the forecasting method currently in use is the correct one. A tracking signal that changes according to the forecast bias shows bias in the prediction model. It is often employed when the forecasting model’s validity is questionable.
@@ 6.3 Phantom inventory
  • The supply of safety may be raised. The enhanced security inventory aims to mitigate inventory problems by having ‘excess’ inventories at hand. RFID may allow the costs of storing this additional and redundant inventory to be reduced.
@@
Inventory precision is a determining factor for forecasting, procurement, and replenishment quality, where inventory records are used as input. Inaccurate demand forecasting due to phantom inventory (overstated PI) may be improved by including RFID in the process [57].
@@ 7 Challenges
This section aims to provide a comprehensive overview of the challenges encountered during the review of 152 articles from 1969 to 2023 in the field of BDA-SCM Framework for Forecasting, with a specific focus on data preprocessing and ML techniques. The challenges identified herein will serve as a valuable resource for future researchers, enabling them to address and overcome these obstacles, ultimately advancing the domain and contributing to its growth and development.
7.0.1 Data Quality and Reliability
One of the critical challenges observed in the reviewed literature is the issue of data quality and reliability. Many studies acknowledged the presence of incomplete, inconsistent, and erroneous data within supply chain datasets. Future research efforts should focus on developing robust data cleansing, integration, and quality assurance techniques to enhance the reliability and accuracy of the forecasting models.
7.0.2 Scalability and Performance
@@
7.0.3 Variety and Complexity of Data Sources
@@
7.0.4 Feature Engineering and Selection
Effective FE and FS play a crucial role in the performance of ML models. However, the surveyed literature lacked comprehensive approaches to identify the most relevant features for forecasting within the supply chain context. Future researchers should investigate advanced FE techniques, including automated FS, dimensionality reduction, and feature representation, to improve forecasting accuracy.
7.0.5 Model Interpretability and Explainability
@@
7.0.6 Real-Time Data Processing and Analysis
SCM requires real-time monitoring and decision-making capabilities. However, the surveyed literature demonstrated a limited exploration of real-time data processing and analysis techniques for forecasting purposes. Future research efforts should concentrate on developing real-time forecasting frameworks that leverage stream processing, online learning, and adaptive algorithms to handle dynamic and time-sensitive supply chain scenarios.
7.0.7 Privacy and Security Concerns
The integration of big data in SCM raises concerns regarding data privacy and security. The surveyed articles paid limited attention to these challenges, and there is a lack of comprehensive approaches to ensure the privacy and security of sensitive supply chain data. Future researchers should focus on developing robust privacy-preserving and secure ML techniques to safeguard data while maintaining the accuracy and efficiency of forecasting models.
7.0.8 Integration of Domain Knowledge
SCM involves complex domain-specific knowledge, including industry-specific constraints, regulations, and contextual factors. The reviewed literature showed a limited integration of such domain knowledge into the forecasting frameworks. Future research should emphasize the incorporation of domain expertise and contextual information to enhance the relevance and accuracy of forecasting models within the supply chain domain.
7.0.9 Lack of Benchmark Datasets and Evaluation Metrics
The absence of standardized benchmark datasets and evaluation metrics hinders the comparison and reproducibility of forecasting techniques. The reviewed articles often utilized different datasets and evaluation metrics, making it challenging to assess the performance of various models. Future researchers should strive to establish benchmark datasets and evaluation protocols specific to supply chain forecasting, enabling fair comparisons and facilitating advancements in the field.
@@ 8 Conclusions
This study has successfully addressed the challenges of SC forecasting and presented cutting-edge technological solutions to overcome them. It has introduced a comprehensive BDA-SCM framework encompassing three essential SC processes: Pre-Process, Control-Process, and Post-Process. The key findings and contributions of this study can be summarized as follows:
@@
    Pre-Process: In the pre-processing stage of SC forecasting, the significance of accurate data aligned with SC objectives was emphasized. The study provided recommendations for SC analysts, including the utilization of exploratory data analysis, FE, hyperparameter tuning, and recent ML model training approaches to improve forecasting accuracy. However, it is essential to note that further research is needed to explore advanced techniques for data cleansing, integration, and quality assurance to ensure reliable and high-quality input data.
@@
    Control-Process: The study discussed how BD could facilitate efficient managerial decision-making in various areas of SCM, such as production and capacity planning, workforce requirements, and inventory management. Leveraging insights from forecasted data allows decision-makers to optimize SC operations and resource allocation. However, future research should focus on developing real-time decision support systems that can integrate and analyze large-scale data streams to enable timely and effective decision-making.
@@
    Post-Process: The post-process section emphasized SC performance measurement and the role of BDA in optimizing model predictions. By analyzing performance metrics and leveraging BDA techniques, SC practitioners can identify areas for improvement and refine their forecasting models accordingly. Future research efforts should focus on developing comprehensive performance measurement frameworks specific to SC forecasting, including quantitative and qualitative metrics, on enabling more accurate evaluation and comparison of forecasting models.
    The accuracy of inventory records is a crucial determinant for forecasting, procurement, and replenishment quality, as these records serve as inputs for these processes. Inaccurate demand forecasting resulting from phantom inventory (overstated PI) can be mitigated through the inclusion of RFID technology in the inventory management process, as evidenced by previous studies. Future research in this area should explore advanced techniques and methodologies for addressing phantom inventory, incorporating emerging technologies, and developing comprehensive frameworks for inventory management and forecasting to enhance the overall performance of supply chain operations.
@@
RQ1: The study has identified and outlined efficient steps to formulate an ML forecasting model for predicting SC factors. Recommendations for accurate data preprocessing, feature engineering, hyperparameter tuning, and advanced ML model training approaches have been provided to enhance the accuracy of SC forecasting models.
@@
RQ3: The study has explored the impact of forecasting on SC performance and identified relevant ML forecasting models for SC forecasting. The connection between accurate forecasting and improved SC performance has been highlighted, with recommendations for performance measurement and the utilization of BDA techniques to optimize model predictions.
By successfully addressing these research questions, this study contributes to the advancement of the field by providing insights into efficient ML modeling steps, the integration of forecasting and SC decision-making processes, and the relevance of ML forecasting models for SC forecasting. Future research should build upon these findings to further enhance the understanding and implementation of BDA in SCM.
This SLR followed a rigorous and objective evaluation approach, utilizing predefined criteria and conducting a comprehensive search to ensure a comprehensive understanding of the BDA-SCM domain. However, it is crucial to acknowledge the limitations of this study. The limitations include the availability of relevant literature, potential publication bias, and the dynamic nature of the BDA-SCM field. Future research should aim to address these gaps by conducting further empirical studies, developing benchmark datasets, and exploring emerging technologies and methodologies to advance the understanding and implementation of BDA in SCM.
By considering and addressing the aforementioned challenges and limitations, future researchers can build upon the findings of this study and contribute to the advancement of the BDA-SCM domain.

### 07 Feb 2024 v5

@@ 5.3 Employee productivity
Data-driven decision-making, or the forecasting of required outputs to estimate the required workforce, is an excellent way to minimize such costs of hiring and laying off by adequately scheduling the workforce [6]. Through proper scheduling, the workforce from idle time can be shifted for workdays requiring extra hours and thereby balanced. [100] showed that work pressure could be balanced with reduced slack time and workforce through different heuristic algorithms, with each algorithm performing well in different areas of efficiency.
@@ 6.1 SC performance
Immediately after the tasks are completed, the performance data must be recorded. For data collection, the performance metrics are first to be identified, just as those found for business evaluations [15]. [51] mentioned plan success, source optimization, production efficiency, delivery performance, and customer support-relation and satisfaction, each having multiple performance metrics under them. The performance level found afterward can be of three types: below average, average, and above-average [5]. The actions followed after such a finding are different in each case. When a below-average performance is observed, managers can either look for anomalies in the system or review whether the goals set were too high to achieve.
@@ 6.2.5 Mean absolute scaled error
With MAE, outliers are protected; with RMSE, we are assured of an impartial prediction. SC Analysts need to analyze MAE and see whether it results in a significant bias; therefore, they should utilize RMSE. In situations when there are many outliers in the dataset, MAE may help correct the skewed prediction.
@@ 6.2.6 Tracking signal
The tracking signal is the way to verify if the current forecasting method is correct. A tracking signal that changes according to the forecast bias shows bias in the prediction model. It is often employed when the forecasting model’s validity is questionable.
@@ 6.3 Phantom inventory
  • The supply of safety may be raised. The enhanced security inventory aims to mitigate inventory problems by having ‘excess’ inventories at hand. RFID may reduce the costs of storing this additional and redundant inventory.
@@
Inventory precision determines forecasting, procurement, and replenishment quality, where inventory records are used as input. Inaccurate demand forecasting due to phantom inventory (overstated PI) may be improved by including RFID in the process [57].
@@ 7 Challenges
This section aims to provide a comprehensive overview of the challenges encountered during the review of 152 articles from 1969 to 2023 in the field of BDA-SCM for forecasting, with a specific focus on data preprocessing and ML techniques. The challenges identified herein will serve as a valuable resource for future researchers, enabling them to address and overcome these obstacles, ultimately advancing the domain and contributing to its growth and development.
7.1 Data quality and reliability
One of the critical challenges observed in the reviewed literature is the issue of data quality and reliability. Many studies acknowledged the presence of incomplete, inconsistent, and erroneous data within SC datasets. Future research efforts should focus on developing robust data cleansing, integration, and quality assurance techniques to enhance the reliability and accuracy of the forecasting models.
7.2 Scalability and performance
@@
7.3 Variety and complexity of data Sources
@@
7.4 Feature Engineering and Selection
Effective FE and to improve forecasting accuracy, including automated FS, dimensionality reduction, and feature representation approaches to identify the most relevant features for forecasting within the SC context. Future researchers should investigate advanced FE techniques, including automated FS, dimensionality reduction, and feature representation, to improve forecasting accuracy.
7.5 Model interpretability and explainability
@@
7.6 Real-time data processing and analysis
SCM requires real-time monitoring and decision-making capabilities. However, the surveyed literature demonstrated a limited exploration of real-time data processing and analysis techniques for forecasting purposes. Future research efforts should concentrate on developing real-time forecasting frameworks that leverage stream processing, online learning, and adaptive algorithms to handle dynamic and time-sensitive SC scenarios.
7.7 Privacy and security concerns
Integrating big data in SCM raises concerns regarding data privacy and security. The surveyed articles paid limited attention to these challenges, and there is a lack of comprehensive approaches to ensure the privacy and security of sensitive SC data. Future researchers should focus on developing robust privacy-preserving and secure ML techniques to safeguard data while maintaining the accuracy and efficiency of forecasting models.
7.8 Integration of domain knowledge
SCM involves complex domain-specific knowledge, including industry-specific constraints, regulations, and contextual factors. The reviewed literature showed a limited integration of such domain knowledge into the forecasting frameworks. Future research should emphasize the incorporation of domain expertise and contextual information to enhance the relevance and accuracy of forecasting models within the SC domain.
7.9 Lack of benchmark datasets and evaluation metrics
The absence of standardized benchmark datasets and evaluation metrics hinders the comparison and reproducibility of forecasting techniques. The reviewed articles often utilized different datasets and evaluation metrics, making it challenging to assess the performance of various models. Future researchers should strive to establish benchmark datasets and evaluation protocols specific to SC forecasting, enabling fair comparisons and facilitating advancements in the field.
@@ 7.9 Lack of benchmark datasets and evaluation metrics
8 Practical implications
The findings of this research offer substantial practical implications for SC practitioners, providing actionable insights that can be effectively implemented in real-world scenarios. The proposed BDA-SCM framework serves as a strategic guide, and its practical application holds the potential for significant benefits in enhancing overall SC operations.
SC practitioners can implement the BDA-SCM framework by initially aligning data collection methodologies with specific SC objectives. This involves adopting a systematic approach to gathering data directly relevant to the SC ecosystem’s unique dynamics and challenges. By integrating the framework into their operational processes, practitioners can leverage the power of BDA at various stages, from problem identification to performance evaluation.
The implementation of the BDA-SCM framework promises several tangible benefits for SC practitioners. Firstly, the framework enhances the accuracy of forecasting models, providing practitioners with more reliable insights into demand patterns, inventory needs, and workforce requirements. This, in turn, enables optimized decision-making across various facets of SCM. Secondly, the cyclic connection within the framework ensures adaptability to dynamic SC conditions. SC practitioners can continuously refine and optimize their forecasting models based on real-time data, thereby staying responsive to changing market dynamics and mitigating potential disruptions. Furthermore, the framework’s emphasis on KPIs and error-measurement systems enables practitioners to evaluate and improve the performance of their forecasting models systematically. This enhances operational transparency and contributes to the overall efficiency and planning effectiveness of the SC.
In practical terms, the BDA-SCM framework supports inventory management by providing accurate demand forecasts, aids in determining workforce needs, optimizes cost factors, and facilitates efficient production and capacity planning. By fostering a holistic approach to SCM, the framework equips practitioners with a systematic and data-driven strategy to address the intricacies of modern SC dynamics. In essence, the practical implementation of the BDA-SCM framework empowers SC practitioners to navigate the complexities of their operational environments with greater precision and foresight, ultimately contributing to enhanced resilience, efficiency, and competitiveness in the ever-evolving SCM landscape.
@@ 9 Conclusions
This systematic review diligently identified and compared state-of-the-art SC forecasting strategies and technologies within the defined temporal scope, conducting a comprehensive review of 152 papers spanning from 1969 to 2023. This study has made significant strides in addressing the challenges inherent in SC forecasting, offering cutting-edge technological solutions within the context of a comprehensive BDA-SCM framework. The key findings and contributions of this study can be summarized as follows:
@@
    Pre-process: In the pre-processing stage of SC forecasting, the significance of accurate data aligned with SC objectives was emphasized. The study provided recommendations for SC analysts, including using EDA, FE, hyperparameter tuning, and recent ML model training approaches to improve forecasting accuracy. However, it is essential to note that further research is needed to explore advanced techniques for data cleansing, integration, and quality assurance to ensure reliable and high-quality input data.
@@
    Control-process: The study discussed how BD could facilitate efficient managerial decision-making in various areas of SCM, such as production & capacity planning, workforce requirements, and inventory management. Leveraging insights from forecasted data allows decision-makers to optimize SC operations and resource allocation. However, future research should focus on developing real-time decision support systems that can integrate and analyze large-scale data streams to enable timely and effective decision-making.
@@
    Post-process: The post-process section emphasized SC performance measurement and the role of BDA in optimizing model predictions. By analyzing performance metrics and leveraging BDA techniques, SC practitioners can identify areas for improvement and refine their forecasting models accordingly. Future research efforts should focus on developing comprehensive performance measurement frameworks specific to SC forecasting, including quantitative and qualitative metrics, to enable more accurate evaluation and comparison of forecasting models. Additionally, the study addresses the accuracy of inventory records as a crucial determinant for forecasting, procurement, and replenishment quality. Mitigating inaccuracies resulting from phantom inventory is highlighted, with the inclusion of RFID technology in inventory management processes as a viable solution. Future research should explore advanced techniques and methodologies to address phantom inventory, incorporating emerging technologies and developing comprehensive inventory management and forecasting frameworks to enhance overall SC performance.
@@
RQ1: The study has identified and outlined efficient steps to formulate an ML forecasting model for predicting SC factors. Recommendations for accurate data preprocessing, FE, hyperparameter tuning, and advanced ML model training approaches have been provided to enhance the accuracy of SC forecasting models.
@@
RQ3: The study has explored the impact of forecasting on SC performance and identified relevant ML forecasting models for SC forecasting. The connection between accurate forecasting and improved SC performance has been highlighted, with recommendations for performance measurement and using BDA techniques to optimize model predictions.
By successfully addressing these RQs, this study contributes to the advancement of the field by providing insights into efficient ML modeling steps, the integration of forecasting and SC decision-making processes, and the relevance of ML forecasting models for SC forecasting. Future research should build upon these findings to further enhance the understanding and implementation of BDA in SCM.
While this systematic literature review (SLR) followed a rigorous and objective evaluation approach, acknowledging its limitations is crucial. These limitations include the availability of relevant literature, potential publication bias, and the dynamic nature of the BDA-SCM field. Future research endeavors should aim to address these gaps by conducting further empirical studies, developing benchmark datasets, and exploring emerging technologies and methodologies to advance the understanding and implementation of BDA in SCM.
By considering and addressing the challenges and limitations outlined in this study, future researchers can build upon its findings and contribute meaningfully to the continued advancement of the BDA-SCM domain.